# SAM-Assisted Remote Sensing Imagery Semantic Segmentation with Object and Boundary Constraints

### 05 Dec 2023 v1

@@ IV-E Ablation Study
To emphasize the distinct roles of the two loss functions in the proposed framework, we conducted ablation experiments using ResNet-18 on two datasets, as detailed in Table III and Table IV. The results underscore that the independent use of these two loss functions enhanced the overall performance of the segmentation model, validating the value and efficacy of SAM’s raw output. When examining the results on impervious surface and building of the ISPRS Vaihingen dataset, it is evident that only the boundary loss exhibited comparable improvement as compared to the combined loss. This suggests that in specific tasks, such as building detection, SAM-Generation results can be fully explored using only the boundary loss. However, in semantic segmentation tasks, where various ground objects possess highly complex boundaries, leveraging both SGO and SGB to their fullest extent becomes necessary. Furthermore, the results of these two loss functions on different categories, as listed in Table III and Table IV, highlight the disparate behaviors exhibited across various categories. By combining the object and boundary loss functions, we proposed a versatile and streamlined framework for directly leveraging SGO and SGB, showcasing robust generalization capabilities in remote sensing image semantic segmentation tasks.
@@ V Conclusion
This work proposed a simple and versatile framework designed to fully exploit the raw output of SAM in conjunction with a segmentation model. Acknowledging the distinctions between remote sensing images and natural images, as well as the characteristics of SGO and SGB, we developed an auxiliary optimization strategy by exploiting two novel loss functions, namely the object loss and the boundary loss. This strategy facilitates improvements in fundamental semantic segmentation tasks without requiring additional network modules. Notably, our introduction of object loss, motivated by the consistency of objects, represents the initial loss function capable of directly utilizing SGO without semantic information. Our validation conducted on two publicly available datasets underscores the robust performance of our framework. Finally, it is worth emphasizing that this work has initiated a preliminary exploration of SAM’s raw output, revealing the potential for large models such as SAM in the field of remote sensing.

### 20 Dec 2023 v2

@@ IV-F Ablation Study
To emphasize the distinct roles of the two loss functions in the proposed framework, we conducted ablation experiments using UNetformer on two datasets, as detailed in Table III and Table IV. The results underscore that the independent use of these two loss functions enhanced the overall performance of the semantic segmentation model, validating the value and efficacy of SAM’s raw output. When examining the results on impervious surface and building of the ISPRS Vaihingen dataset, it is evident that only the boundary preservation loss exhibited comparable improvement as compared to the combined loss. This suggests that SAM-based pre-processing results can be fully explored using only the boundary preservation loss in specific tasks, such as building detection. However, in semantic segmentation tasks, where various ground objects possess highly complex boundaries, leveraging both SGO and SGB to their fullest extent becomes necessary. Furthermore, the results of these two loss functions on different categories, as listed in Table III and Table IV, highlight the disparate behaviors exhibited across various categories. Combining the object and boundary preservation loss functions, we proposed a versatile and streamlined framework for directly leveraging SGO and SGB, showcasing robust generalization capabilities in remote sensing image semantic segmentation tasks.
IV-G Model Complexity Analysis
We evaluate the computational complexity of the proposed framework using various metrics: floating point operation count (FLOPs), model parameter, memory footprint, running time (s) and frames per second (FPS). FLOPs accesses the model’s complexity, while model parameters and memory footprint evaluate the scale of the network and memory requirements, respectively. Finally, running time and FPS quantifies and training speed and inference speed, respectively. Ideally, an efficient model maintains lower values in FLOPs, model parameters, memory footprint and running time while achieving a higher FPS value.
Table V shows the complexity analysis results of all comparing semantic segmentation models considered in this work. Inspection of Table V shows that our approach introduces no extra model complexity or inference time. In our framework, it is necessary to generate SGO and SGB before their utilization in loss calculation, hereby negating the requirement for additional task-specific modules. However, it is observed that our approach lengthens the training time since the gradient back-propagation takes longer to compute with two additional loss functions. On the other hand, during inference stage as presented in Fig. 3 (b), the proposed approach operates identically to the original model, ensuring zero impact on inference speed. Considering the enhancements observed in other aspects, we believe that the marginal rise in training time is a justifiable trade-off. These outcomes highlight the remarkable scalability and wide application of our framework across current semantic segmentation models.
@@ V Conclusion
This work proposed a simple and versatile framework designed to fully exploit the raw output of SAM in conjunction with general remote sensing imagery semantic segmentation models. Acknowledging the distinctions between remote sensing images and natural images, as well as the characteristics of SGO and SGB, we developed an auxiliary optimization strategy by exploiting two loss functions, object consistency loss and boundary preservation loss. This strategy facilitates improvements in fundamental semantic segmentation tasks with different network structures without requiring additional modules. Notably, our introduction of object consistency loss, motivated by the consistency of objects, represents the initial loss function capable of directly utilizing SGO without semantic information. Our validation on two publicly available datasets and four comprehensive semantic segmentation models highlights the robust performance of our framework. Finally, it is worth emphasizing that this work has initiated a preliminary exploration of SAM’s raw output, revealing the potential for large models such as SAM in remote sensing.